# Hierarchical modeling of gravitational-wave populations for disentangling environmental and modified-gravity effects

### v1

@@ VI.1 Summary
We check that the Bayes factor $\mathcal{B}^{g}_{\ell}$ scales proportionally to the intrinsic strength $\dot{G}^{*}[\rm yr]^{-1}$ of the global effect in the population, and inversely with the fraction of sources affected by the local effect, $f^{*}$. We also find that the expectation of the global effect hyperparameter $\langle\dot{G}\rangle$ may significantly deviate from the ground truth if $f^{*}\geq 0.9$. We further highlighted a scenario where the Bayes factor $\log_{10}\mathcal{B}^{g}_{\ell}\approx 0.0$ even when both effects are present in the population, such that the ability to tell them apart with the data is lost. Future studies should study such situations in more detail to further assess the robustness of the hierarchical framework.
@@ VI.2 Future directions and outlook
While we only consider circular-equatorial Kerr EMRIs in this paper, future studies should study the robustness of the hierarchical framework for beyond-vacuum-GR effects in more generic orbits, i.e., including the effects of eccentricity and inclination in the data. Such extensions may reduce parameter space degeneracies and, therefore, the biases incurred in recovering the local and global effect hyperparameters in mixed populations. More accurate models for EMRI waveforms and beyond-vacuum-GR effects are becoming increasingly available Chapman-Bird et al. (2025); Cardoso et al. (2022); Duque et al. (2025a); Zi and Kumar (2025); Duque et al. (2025b).
@@
Finally, we caution that some systematic biases may persist in the hierarchical framework, e.g., due to inaccuracies in waveform modeling which can mimic global effects, as discussed in Moore et al. (2021), and should be addressed in future work. In our framework, these systematic biases can be understood to be induced by “local” waveform modeling inaccuracies, which may strongly depend on the source parameters Khalvati et al. (2025), biasing the inference of the intrinsic strength $\dot{G}$ of the global effect. Our setup can be readily employed to study such systematics in the future. Extensions of the current framework, shaped by these careful considerations, will pave the way for robust and realistic tests of GR in the presence of astrophysical environments with the next generation of GW detectors.
@@ Acknowledgements.
SK thanks Patrick Meyers and Michele Vallisneri for useful discussions during a research visit at ETH Zurich. SK also acknowledges the computing resources provided by the NUS IT Research Computing group and the support of the NUS Research Scholarship. EB acknowledges support from the European Union’s Horizon ERC Synergy Grant “Making Sense of the Unexpected in the Gravitational-Wave Sky” (Grant No. GWSky-101167314) and the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”. This work has been supported by the Agenzia Spaziale Italiana (ASI), Project n. 2024-36-HH.0, “Attività per la fase B2/C della missione LISA”.

### v2

@@ VI.1 Summary
In addition, for two different populations of sizes $N_{\rm pop}=100,500$, we check that the Bayes factor $\mathcal{B}^{g}_{\ell}$ scales proportionally to the intrinsic strength $\dot{G}^{*}[\rm yr]^{-1}$ of the global effect in the population, and inversely with the fraction of sources affected by the local effect, $f^{*}$. We also find that the expectation of the global effect hyperparameter $\langle\dot{G}\rangle$ may significantly deviate from the ground truth if $f^{*}\geq 0.8$ when $N_{\rm pop}=500$. For both cases $N_{\rm pop}=100,500$, we highlighted scenarios where the Bayes factor $\log_{10}\mathcal{B}^{g}_{\ell}\approx 0.0$ even when both effects are present in the population, such that the ability to tell them apart with the data is lost. Future studies should study such situations in more detail to further assess the robustness of the hierarchical framework.
@@ VI.2 Future directions and outlook
While we only consider circular-equatorial Kerr EMRIs in this paper, future studies should study the robustness of the hierarchical framework for beyond-vacuum-GR effects in more generic orbits, i.e., including the effects of eccentricity and inclination in the data. Such extensions may reduce parameter space degeneracies and, therefore, the biases incurred in recovering the local and global effect hyperparameters in mixed populations. More accurate models for EMRI waveforms and beyond-vacuum-GR effects are becoming increasingly available Chapman-Bird and others (2025); Cardoso et al. (2022); Duque et al. (2025a); Zi and Kumar (2025); Duque et al. (2025b); Hegade K. R. et al. (2025b, a).
@@
Finally, we caution that some systematic biases may persist in the hierarchical framework, e.g., due to inaccuracies in waveform modeling which can mimic global deviations from the null hypothesis, as discussed in Moore et al. (2021). In our framework, these systematic biases can be understood to be induced by modeling errors, such as missing higher-order corrections in the vacuum-GR evolution Burke et al. (2024) or insufficiently accurate template waveforms Khalvati et al. (2025). Because such waveform systematics will be pervasive throughout the observed population, they may strongly bias the inference of a global effect such as the intrinsic strength $\dot{G}$, even within the hierarchical framework. Our setup can be readily employed to study such systematics in the future. Extensions of the current framework, shaped by these careful considerations, will pave the way for robust and realistic tests of GR in the presence of astrophysical environments with the next generation of GW detectors.
@@ Acknowledgements.
SK thanks Patrick Meyers and Michele Vallisneri for useful discussions during a research visit to ETH Zurich. SK also acknowledges the computing resources provided by the NUS IT Research Computing group and the support of the NUS Research Scholarship. EB acknowledges support from the European Union’s Horizon ERC Synergy Grant “Making Sense of the Unexpected in the Gravitational-Wave Sky” (Grant No. GWSky-101167314) and the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”. This work has been supported by the Agenzia Spaziale Italiana (ASI), Project n. 2024-36-HH.0, “Attività per la fase B2/C della missione LISA”.